For your science exhibit, you want to make a block of wood float at the interface between a

layer of oil of unknown density 𝜌oil, and water (π†π°πšπ­πžπ«=𝟏𝟎𝟎𝟎 [kg/m3]) inside a container as shown. The density of the wood is 𝝆𝐰𝐨𝐨𝐝=πŸ–πŸŽπŸŽ [π€π π¦πŸ‘β„]. What must be the density of the oil so that only πŸ‘πŸŽ.𝟎% of the wood is submerged in water?



1
Expert's answer
2021-06-01T10:50:54-0400

A fraction k of block volume must be submerged in water.

The force of gravity acting on the wood is


"F=mg=\rho gV."

The buoyancy force in water:


"F_1=\rho_1 g·kV."

The buoyancy force in oil:


"F_2=\rho g·(1-k)V."

The two buoyancy forces prevent the block from sinking:


"F_1+F_2=F,\\\space\\
k\rho_1+(1-k)\rho_2=\rho,\\\space\\
\rho_2=\frac{\rho-k\rho_1}{1-k}=714.29\text{ kg/m}^3."